THE VLA EXPANSION PROJECT

A Presentation to the Radio/Submm Panel of the

Astronomy Decadal Review February 19, 1999

Rick Perley

National Radio Astronomy Observatory

Project Goal

• To improve, typically by a factor of ten, the technical capability of the Very Large Array in every key observational characteristic:

– Senstivity

– Frequency Accessibility

– Spectral Capability

– Resolution

– Surface Brightness Sensitivity

Possible observations include

• 3D solar tomography (cyclotron lines at diff. depths)

• imaging the dynamic heliosphere (through scattering & RM)

• census of stellar activity in normal MS stars • Galactic center (orientations and strengths

from 0.1 to 100s of pc)

• galaxy haloes (spectral index mapping, pol’n)

• galaxy clusters (see next page)

• intergalactic mag. fields (Faraday rotation along adjacent lines-of-sight)

• high-redshift sources (both low and high RM; synchrotron mapping)

Galaxy Clusters with the Expanded VLA